For the pair of functions, find f(g(x)) and g(f(x)).
F(x)=3x, g(x)=x squared

Answer by stanbon(75887)

find f(g(x)) and g(f(x)).
f(x)=3x,
g(x)=x^2
---------------
f(g(x)) = f(x^2) = 3x^2
-----------
g(f(x)) = g(3x) = (3x)^2 = 9x^2
=======================================
Cheers,
Stan H.